# VADB: A Large-Scale Video Aesthetic Database with Professional and Multi-Dimensional Annotations

### v1

@@ 1 Introduction
To address this, the development of a scientific and systematic video aesthetic evaluation framework, the creation of a high-quality, multidimensional video aesthetic dataset with rich annotations, and the design of robust, generalizable, and accurate assessment models are critical research objectives. Unlike image aesthetic assessment, which leverages established principles like the golden ratio and color harmony (Yi Lu, 2021)(Houlgate, 2009), video aesthetic assessment is more complex due to its dynamic spatiotemporal nature. A comprehensive evaluation system thus requires thorough investigation of key aesthetic attributes shaping viewer perception and the establishment of tailored standards for diverse video genres. Methodologically, integrating objective technical metrics—such as camera movement and composition types—with traditional subjective evaluations is essential for a balanced assessment approach.
@@
While image aesthetic assessment has advanced significantly in areas like classification, score distribution, and attribute analysis (Jin et al., 2023)(Jin et al., 2024a)(Huang et al., 2024), video aesthetic assessment lags due to technical challenges and the scarcity of large-scale datasets. Notably, the pretrained CLIP model(Radford et al., 2021) has shown exceptional performance in image aesthetic assessment(Sheng et al., 2023)(Jin et al., 2024b) and video retrieval(Luo et al., 2021). Utilizing multimodal video aesthetic datasets with detailed language annotations for CLIP-based pretraining offers substantial potential to overcome current limitations in video aesthetic research.
@@ 2.1 Video Aesthetic Dataset
To highlight the limitations of video aesthetic datasets, we first review the advancements in video quality datasets as a benchmark. Video quality datasets have progressed significantly, evolving from laboratory-controlled compression artifacts (Wang et al., 2016, 2017) to real-world scenarios with mixed distortions (Hosu et al., 2017; Sinno and Bovik, 2019). Annotation methodologies have advanced from single-score ratings (Sinno and Bovik, 2019) to multi-dimensional quality attributes and textual descriptions (Hosu et al., 2017; Duan et al., 2024). Concurrently, dataset scales have increased from hundreds of clips (Li et al., 2020) to tens of thousands, with annotations reaching millions (Ying et al., 2021; Duan et al., 2024).
By contrast, video aesthetic quality datasets have developed more slowly, constrained by challenges in annotation standardization and high labeling costs. Early efforts, such as the Telefonica dataset (Bylinskii et al., 2015) with 160 YouTube videos annotated using five-level aesthetic ratings, established initial benchmarks. Subsequent datasets introduced greater diversity, including Niu et al.’s 2,000 professional and amateur videos with binary classification (Niu and Liu, 2012) and the VAQ700 dataset with 700 daily-life videos and multi-annotator ratings (Tzelepis et al., 2016). Larger-scale datasets, such as AVAQ6000 (Kuang et al., 2019) with 6,000 drone videos, relied on binary labels, limiting their annotation richness. More recently, DIVIDE-3k (Wu et al., 2023) provided multi-dimensional ratings for 3,590 videos. Despite these advances, video aesthetic datasets remain constrained by limited scale, inconsistent annotations, and insufficient dimensional richness compared to video quality datasets, indicating a need for more comprehensive solutions.
@@ 2.2 Video Aesthetic Assessment Methods
Traditional methods rely on photographic principles, using handcrafted features to evaluate visual appeal.  (Luo and Tang, 2008) extracted subject regions and motion features to distinguish professional from amateur videos.  (Moorthy et al., 2010) utilized features like motion amplitude ratio and adherence to the rule of thirds, pooling frame-level features into video-level representations.  (Niu and Liu, 2012) combined static image aesthetic features with dynamic features, for multi-class quality classification in professional video production.  (Yeh et al., 2013) introduced motion features derived from optical flow within a temporal-aware framework.  (Peng et al., 2021) evaluated robotic dance aesthetics by integrating spatial and shape features. These methods, while effective, are limited in capturing complex dynamic patterns.
Deep learning approaches, though constrained by limited video aesthetic datasets, show significant promise.  (Phatak et al., 2019) applied CNNs to assess the impact of motion speed on aesthetics, outperforming handcrafted features. (Kuang et al., 2020) proposed a multimodal CNN that fused drone video, trajectory, and 3D structural features for aesthetic classification. (Asarkar and Phatak, 2019) integrated color contrast features with deep learning to explore connections between video emotion and aesthetics. (Wu et al., 2023) introduced the DOVER video quality assessor, combining aesthetic and technical perspectives to evaluate user-generated content (UGC) video quality. Despite fewer studies, these advancements highlight deep learning’s potential to enhance aesthetic assessment.
@@ 2.3 Learning Visual Representations from Text Supervision
Text-supervised visual representation learning, leveraging abundant internet image-text pairs, has become a prominent research area. CLIP (Radford et al., 2021) learns robust visual representations through contrastive language-image pre-training, enabling effective transfer to tasks like image retrieval (Baldrati et al., 2023) and aesthetic assessment (Sheng et al., 2023; Jin et al., 2024b). In video research, ClipBERT (Lei et al., 2021) adapts image-text pre-training for video question answering and text-to-video retrieval, while Clip4Clip (Luo et al., 2021) and X-CLIP (Ni et al., 2022) extend this approach to video-text pre-training, improving video retrieval performance. Inspired by these approaches, VADB-Net applies CLIP’s representations to video aesthetic assessment.
@@ 3.1 Video Categories & Attributes & Tags
The attributes listed in Table 1 are grounded in the aesthetic framework of film and television studies (Wu, 2024; Matbouly, 2022; Petrogianni et al., 2022; Arijon, 2013), enabling a comprehensive evaluation of a video’s aesthetic quality. Notably, given the dynamic nature of subjects and the richness of emotional expression in character videos, specific character-specific attributes are annotated in addition to general attributes.
@@ 4.1 Pre-training Stage
The Video Encoder, built upon CLIP ViT-B/32(Radford et al., 2021), employs 3D convolution initialization to extend 2D convolutional kernels for extracting spatiotemporal features. After 12-layer Transformer encoding, frame-level mean pooling generates fixed-length video representations.
@@ 4.3.2 Comparative Experiment
Although video aesthetic scoring models are scarce in academia, recent progress in video quality assessment has been notable. The FAST-VQA framework (Wu et al., 2022) efficiently learns quality-related features through end-to-end training, while SimpleVQA (Sun et al., 2022) effectively extracts spatial and temporal features. Furthermore, ModularBVQA (Wen et al., 2024) achieves accurate predictions by jointly processing visual content, resolution, and frame rate. Table 4 presents the comparative experimental results between VADB-Net and these three models.
@@ Appendix C Selection of Video Aesthetic Attributes
To construct a broadly applicable dataset for video aesthetic scoring and commentary, this study meticulously selects and designs both universal and specialized aesthetic attributes based on film and media aesthetics theory (Wu, 2024; Matbouly, 2022; Petrogianni et al., 2022; Arijon, 2013). These attributes comprehensively evaluate the visual expressiveness of videos, accommodating the diverse sources and varying quality levels of the dataset while providing a unified standard for assessing different video types.
@@ E.1.1 Model Architecture
Refer to CLIP4CLIP(Luo et al., 2021), the Pre-training Stage model extends the CLIP framework (ViT-B/32) for video-text retrieval, comprising a video encoder, dual text encoders, and a dynamic fusion module. The video encoder leverages CLIP’s 12-layer visual Transformer (vision_layers=12), employing 3D convolution initialization (linear_patch="3d") to extend 2D convolutional kernels for capturing spatiotemporal features. It processes 12 video frames (max_frames=12, resolution $\times224$) and generates a 512-dimensional video representation via mean pooling (sim_header="meanP").

### v2

@@ 1 Introduction
To address this, the development of a scientific and systematic video aesthetic evaluation framework, the creation of a high-quality, multidimensional video aesthetic dataset with rich annotations, and the design of robust, generalizable, and accurate assessment models are critical research objectives. Unlike image aesthetic assessment, which leverages established principles like the golden ratio and color harmony (Fundamentalsofmold-making)(sep-hegel-aesthetics), video aesthetic assessment is more complex due to its dynamic spatiotemporal nature. A comprehensive evaluation system thus requires thorough investigation of key aesthetic attributes shaping viewer perception and the establishment of tailored standards for diverse video genres. Methodologically, integrating objective technical metrics—such as camera movement and composition types—with traditional subjective evaluations is essential for a balanced assessment approach.
@@
While image aesthetic assessment has advanced significantly in areas like classification, score distribution, and attribute analysis (jin2023aesthetic)(jin2024paintings)(huang2024aesexpert), video aesthetic assessment lags due to technical challenges and the scarcity of large-scale datasets. Notably, the pretrained CLIP model(radford2021learning) has shown exceptional performance in image aesthetic assessment(sheng2023aesclip)(jin2025paintings) and video retrieval(luo2021clip4clip). Utilizing multimodal video aesthetic datasets with detailed language annotations for CLIP-based pretraining offers substantial potential to overcome current limitations in video aesthetic research.
@@ 2.1 Video Aesthetic Dataset
To highlight the limitations of video aesthetic datasets, we first review the advancements in video quality datasets as a benchmark. Video quality datasets have progressed significantly, evolving from laboratory-controlled compression artifacts (7532610; WANG2017292) to real-world scenarios with mixed distortions (7965673; 8463581). Annotation methodologies have advanced from single-score ratings (8463581) to multi-dimensional quality attributes and textual descriptions (7965673; duan2024finevq). Concurrently, dataset scales have increased from hundreds of clips (9175514) to tens of thousands, with annotations reaching millions (9577503; duan2024finevq).
By contrast, video aesthetic quality datasets have developed more slowly, constrained by challenges in annotation standardization and high labeling costs. Early efforts, such as the Telefonica dataset (BYLINSKII2015258) with 160 YouTube videos annotated using five-level aesthetic ratings, established initial benchmarks. Subsequent datasets introduced greater diversity, including Niu et al.’s 2,000 professional and amateur videos with binary classification (6162974) and the VAQ700 dataset with 700 daily-life videos and multi-annotator ratings (tzelepis2016video). Larger-scale datasets, such as AVAQ6000 (kuang2019deep) with 6,000 drone videos, relied on binary labels, limiting their annotation richness. More recently, DIVIDE-3k (wu2023exploring) provided multi-dimensional ratings for 3,590 videos. Despite these advances, video aesthetic datasets remain constrained by limited scale, inconsistent annotations, and insufficient dimensional richness compared to video quality datasets, indicating a need for more comprehensive solutions.
@@ 2.2 Video Aesthetic Assessment Methods
Traditional methods rely on photographic principles, using handcrafted features to evaluate visual appeal.  (luo2008photo) extracted subject regions and motion features to distinguish professional from amateur videos.  (moorthy2010towards) utilized features like motion amplitude ratio and adherence to the rule of thirds, pooling frame-level features into video-level representations.  (6162974) combined static image aesthetic features with dynamic features, for multi-class quality classification in professional video production.  (yeh2013video) introduced motion features derived from optical flow within a temporal-aware framework.  (peng2021multiple) evaluated robotic dance aesthetics by integrating spatial and shape features. These methods, while effective, are limited in capturing complex dynamic patterns.
Deep learning approaches, though constrained by limited video aesthetic datasets, show significant promise.  (phatak2019deep) applied CNNs to assess the impact of motion speed on aesthetics, outperforming handcrafted features. (8936378) proposed a multimodal CNN that fused drone video, trajectory, and 3D structural features for aesthetic classification. (asarkar2019detecting) integrated color contrast features with deep learning to explore connections between video emotion and aesthetics. (wu2023exploring) introduced the DOVER video quality assessor, combining aesthetic and technical perspectives to evaluate user-generated content (UGC) video quality. Despite fewer studies, these advancements highlight deep learning’s potential to enhance aesthetic assessment.
@@ 2.3 Learning Visual Representations from Text Supervision
Text-supervised visual representation learning, leveraging abundant internet image-text pairs, has become a prominent research area. CLIP (radford2021learning) learns robust visual representations through contrastive language-image pre-training, enabling effective transfer to tasks like image retrieval (10.1145/3617597) and aesthetic assessment (sheng2023aesclip; jin2025paintings). In video research, ClipBERT (lei2021less) adapts image-text pre-training for video question answering and text-to-video retrieval, while Clip4Clip (luo2021clip4clip) and X-CLIP (ni2022expanding) extend this approach to video-text pre-training, improving video retrieval performance. Inspired by these approaches, VADB-Net applies CLIP’s representations to video aesthetic assessment.
@@ 3.1 Video Categories & Attributes & Tags
The attributes listed in Table 1 are grounded in the aesthetic framework of film and television studies (Wu2024; Matbouly2022; Petrogianni2022; arijon2013), enabling a comprehensive evaluation of a video’s aesthetic quality. Notably, given the dynamic nature of subjects and the richness of emotional expression in character videos, specific character-specific attributes are annotated in addition to general attributes.
@@ 4.1 Pre-training Stage
The Video Encoder, built upon CLIP ViT-B/32(radford2021learning), employs 3D convolution initialization to extend 2D convolutional kernels for extracting spatiotemporal features. After 12-layer Transformer encoding, frame-level mean pooling generates fixed-length video representations.
@@ 4.3.2 Comparative Experiment
Although video aesthetic scoring models are scarce in academia, recent progress in video quality assessment has been notable. The FAST-VQA framework (wu2022fast) efficiently learns quality-related features through end-to-end training, while SimpleVQA (sun2022deep) effectively extracts spatial and temporal features. Furthermore, ModularBVQA (wen2024modular) achieves accurate predictions by jointly processing visual content, resolution, and frame rate. Table 4 presents the comparative experimental results between VADB-Net and these three models.
@@ Appendix C Selection of Video Aesthetic Attributes
To construct a broadly applicable dataset for video aesthetic scoring and commentary, this study meticulously selects and designs both universal and specialized aesthetic attributes based on film and media aesthetics theory (Wu2024; Matbouly2022; Petrogianni2022; arijon2013). These attributes comprehensively evaluate the visual expressiveness of videos, accommodating the diverse sources and varying quality levels of the dataset while providing a unified standard for assessing different video types.
@@ E.1.1 Model Architecture
Refer to CLIP4CLIP(luo2021clip4clip), the Pre-training Stage model extends the CLIP framework (ViT-B/32) for video-text retrieval, comprising a video encoder, dual text encoders, and a dynamic fusion module. The video encoder leverages CLIP’s 12-layer visual Transformer (vision_layers=12), employing 3D convolution initialization (linear_patch="3d") to extend 2D convolutional kernels for capturing spatiotemporal features. It processes 12 video frames (max_frames=12, resolution $\times224$) and generates a 512-dimensional video representation via mean pooling (sim_header="meanP").